The interior angles of a polygon are in A.P. if the smallest angle is 120° and common difference is 5° then the number of sides in polygon is ?
Answer by jim_thompson5910(35256)

first term: 120
second term: 120+5
third term: 120+5*2
...
...
...
nth term: 120+5*(n-1)
-------------------------------------------------------
Sum of the arithmetic terms


Sn = n*(a1 + an)/2


Sn = n*(120 + 120+5*(n-1))/2


Sn = n*(120 + 120+5n-5)/2


Sn = n*(5n+235)/2

-------------------------------------------------------

Sum of angles in a polygon


S = 180(n-2)


n*(5n+235)/2 = 180(n-2) ... replace S with the arithmetic sum found above


n*(5n+235) = 2*180(n-2)


5n^2+235n = 360n-720


5n^2+235n - 360n+720 = 0


5n^2-125n+720 = 0


Now use the quadratic formula to solve for n. I'll let you do this part.


If you get any n values that are negative or not whole numbers, then ignore them. The value of n is going to be some positive whole number.